OpenStudy (anonymous):

What does it mean to write f(x)=ln(x+1/3x+2), nontrivially as the composite of two functions?

OpenStudy (anonymous):

let g(x)= x+1/3x+2 , f(x)= ln(x) . hence fog(x)

OpenStudy (anonymous):

hope that helps

OpenStudy (anonymous):

It does, thanks a lot

OpenStudy (anonymous):

fan me pls if u like the answer

OpenStudy (anonymous):

cheers